# The Binary and Ternary Quantization Can Improve Feature Discrimination

### v1

@@ Abstract
In machine learning, quantization is widely used to simplify data representation and facilitate algorithm deployment on hardware. Given the fundamental role of classification in machine learning, it is crucial to investigate the impact of quantization on classification. Current research primarily focuses on quantization errors, operating under the premise that higher quantization errors generally result in lower classification performance. However, this premise lacks a solid theoretical foundation and often contradicts empirical findings. For instance, certain extremely low bit-width quantization methods, such as $\{0,1\}$-binary quantization and $\{0,\pm 1\}$-ternary quantization, can achieve comparable or even superior classification accuracy compared to the original non-quantized data, despite exhibiting high quantization errors. To more accurately evaluate classification performance, we propose to directly investigate the feature discrimination of quantized data, instead of analyzing its quantization error. Interestingly, it is found that both binary and ternary quantization methods can improve, rather than degrade, the feature discrimination of the original data. This remarkable performance is validated through classification experiments across various data types, including images, speech, and texts.
@@ 1 Introduction
For instance, it has been observed that some extremely low bit-width quantization methods, such as the $\{0,1\}$-binary quantization and $\{0,\pm 1\}$-ternary quantization, which have been successively applied in large-scale retrieval (Charikar, 2002) and deep network quantization (Qin et al., 2020; Gholami et al., 2022), can achieve comparable or even superior classification performance than their full-precision counterparts (Courbariaux et al., 2015; Lin et al., 2016b; Lu et al., 2023), despite suffering from high quantization errors. Apparently, the remarkable classification improvement resulting from quantization should not be attributed to the high quantization errors. This reveals the inadequacy of quantization errors in evaluating the actual classification performance. Due to the absence of a theoretical explanation, the classification improvement induced by quantization has often been regarded as incidental and received little attention. Instead of quantization errors, in the paper we demonstrate that this intriguing phenomenon can be reasonably explained by feature discrimination. Following the Fisher’s linear discriminant analysis (Fisher, 1936), we here refer to feature discrimination as the ratio between inter-class and intra-class scatters, and evaluate the classification performance based on the rule that the higher the feature discrimination, the easier the classification. To the best of our knowledge, this is the first study that exploits feature discrimination to analyze the impact of quantization on classification, although it is more direct and reasonable than quantization errors in evaluating classification performance. The scarcity of relevant research can be attributed to the nonlinearity of the quantization operation, which substantially increases the analytical complexity of feature discrimination functions.
In the paper, it is demonstrated that the impact of the threshold-based binary and ternary quantization on feature discrimination can be analyzed, when the data are appropriately modeled using a Gaussian mixture model, with each Gaussian element representing one class of data. The Gaussian mixture model is chosen here based on two considerations. Firstly, the model has been well-established for approximating the distributions of real-world data (Torralba and Oliva, 2003; Weiss and Freeman, 2007) and their feature transformations (Wainwright and Simoncelli, 1999; Lam and Goodman, 2000). Secondly, the closure property of Gaussian distributions under linear operations can simplify the analysis of the feature discrimination function. By analyzing the discrimination across varying quantization thresholds, it is found that there exist certain quantization thresholds that can improve the discrimination of original data, thereby yielding improved classification performance. This finding is extensively validated through classification experiments both on synthetic and real data.
@@ 2.1 Quantization functions
where the threshold parameter $\tau\in(-\infty,+\infty)$ for $f_{b}(x;\tau)$, and $\tau\in[0,+\infty)$ for $f_{t}(x;\tau)$. The two functions operate element-wise on a vector $\mathbf{x}=[x_{1},x_{2},\cdots,x_{n}]^{\top}\in\mathbb{R}^{n}$, namely $f_{b}(\mathbf{x};\tau)=(f_{b}(x_{1};\tau),f_{b}(x_{2};\tau),\cdots,f_{b}(x_{n};\tau))^{\top}$ and the same applies to $f_{t}(\mathbf{x};\tau)$.
@@ 2.2 Data distributions
When standardization, a common practice in machine learning, is applied to the two variables, $X$ and $Y$, their distributions will exhibit specific relationships. More precisely, in a binary classification problem, the dataset we handle is a mixture, denoted as $Z$, comprising two classes of samples drawn respectively from $X$ and $Y$. Usually, the mixture $Z$ is assumed to possess a balanced class distribution, meaning that samples are drawn from $X$ and $Y$ with equal probabilities. Under this assumption, when we perform standardization by subtracting the mean and dividing by the standard deviation for each sample in $Z$, the distributions of $X$ and $Y$ (in $Z$) will become
@@
From Equations (3) and (4), it can be seen that after standardization, the two classes of variables $\tilde{X}$ and $\tilde{Y}$ still exhibit Gaussian distributions, but showcase two interesting properties: 1) their means are symmetric about zero; and 2) they have the sum of the square of the mean and the variance equal to one. In the paper, we will focus our study on such standardized data, with the aforementioned properties specified in Property 1.
Property 1 (The distributions of two classes of standardized data)
The two classes of standardized data we aim to study have their samples i.i.d drawn from $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$, where $\mu^{2}+\sigma^{2}=1$, $\mu\in(0,1)$.
@@ 2.4 Goal
The major goal of the paper is to investigate whether there exist threshold values $\tau$ in the binary quantization $f_{b}(x;\tau)$ and the ternary quantization $f_{t}(x;\tau)$, such that the quantization can improve the feature discrimination of original data, namely having $D_{b}>D$ and $D_{t}>D$.
@@ 3.1 Theoretical results
Remarks: Regarding the two theorems, there are three issues worth discussing. 1) The two theorems suggest that both binary and ternary quantization methods can indeed improve the classification performance of original data, if there exist quantization thresholds $\tau$ that can satisfy the constraints shown in Equations (8) and (9). The following numerical analysis demonstrates that the desired threshold $\tau$ does exist, when the two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ are assigned appropriate values for $\mu$ and $\sigma$. This threshold $\tau$ can be typically determined using gradient descent methods, as detailed in Appendix B. 2) Our theoretical analysis is based on the premise that the data vectors within the same class follow Gaussian distributions in each dimension of the vectors. This condition should hold true when two classes of data are readily separable, as in this case the data points within each class should cluster tightly, allowing for Gaussian approximation. This explains the recent research findings, that the binary or ternary quantization tends to achieve comparable or superior classification performance, when handling relatively simple datasets (Courbariaux et al., 2015; Lin et al., 2016b), or distinguishable features (Lu et al., 2023). 3) The conclusion we derive in Theorem 4 for $\{0,1\}$-binary quantization also applies to another popular $\{-1,1\}$-binary quantization (Qin et al., 2020), since the Euclidean distance of the former is equivalent to the cosine distance of the latter.
@@ 3.2 Numerical validation
In this part, we conduct numerical analyses for two primary objectives. Firstly, we aim to prove the existence of the desired quantization threshold $\tau$ that holds Equations (8) and (9), namely making the left sides of the two inequalities larger than their right sides (with values equal to zero). For this purpose, we compute the values of the left sides of Equations (8) and (9), through assigning specific values to $\tau$, as well as to the two variables $X$ and $Y$’s distribution parameters $\mu$ and $\sigma^{2}$. Note that we here set $\sigma^{2}=1-\mu^{2}$, $\mu\in(0,1)$, in accordance with Property 1. In Figure 1, we examine the case that fixes $\mu=0.8$ and $\sigma^{2}=0.36$, while varying the value of $\tau$ with a step width 0.01. The results for binary quantization and ternary quantization are provided in Figures 1 (a) and (c), respectively. It can be seen that for the two quantization methods, their conditions shown in Equations (8) and (9) will hold when respectively having $\tau\in[-0.2,0.2]$ and $\tau\in[0,0.5]$. This proves the existence of the desired quantization threshold $\tau$ that can improve feature discrimination. For limited space, we here only discuss the case of $\mu=0.8$ (and $\sigma^{2}=1-\mu^{2}$) in Figure 1. By examining different $\mu\in(0,1)$ in the same way, we can find that the quantization threshold $\tau$ that holds Equations (8) and (9), is present when $\mu\in(0.76,1)$ and $\mu\in(0.66,1)$, respectively; see Figures 7 and 8 for more evidences. This result implies two consequences. On one hand, ternary quantization has more chances to improve feature discrimination compared to binary quantization, as the former has a broader range of $\mu$. On the other hand, the improved discrimination tends to be achieved when $\mu$ is sufficiently large, coupled with a correspondingly small $\sigma$, or when the discrimination between two classes of data is sufficiently high.
@@
The second goal is to verify that the quantization thresholds $\tau$ we estimate with Equations (8) and (9) in Theorems 4 and 5, can indeed improve feature discrimination. To this end, it needs to prove that the ranges of $\tau$ derived by Equations (8) and (9), such as the ones depicted in Figures 1 (a) and (c), are consistent with the ranges we can statistically estimate by the discrimination definitions $D$, $D_{b}$ and $D_{t}$, as specified in Definitions 2 and 3. To estimate the discrimination $D$, $D_{b}$ and $D_{t}$, we randomly generate 10,000 samples from $X\sim N(0.8,0.36)$ and $Y\sim N(-0.8,0.36)$, respectively, and then statistically estimate them with Equations (5), (6) and (7), across varying $\tau$ (with a step width 0.01). The results are provided in Figures 1 (b) and (d), respectively for binary and ternary quantization. It can be seen that we have $\tau\in[-0.2,0.2]$ for $D_{b}>D$, and have $\tau\in[0,0.5]$ for $D_{t}>D$. The results coincide with the theoretical results shown in Figures 1 (a) and (c), validating Theorems 4 and 5.
@@ 4 Experiments
Through previous theoretical and numerical analyses, we have demonstrated that binary and ternary quantization can improve feature discrimination between two classes of data, when the data vectors within each class exhibit Gaussian distributions across each dimension of their feature vectors. As improved feature discrimination is expected to enhance classification performance, this section aims to confirm this improvement by evaluating classification performance. To see the generalizability of our theoretical findings, we evaluate classification performance on both synthetic and real data. Synthetic data can accurately adhere to the distribution conditions specified in our theoretical analysis, whereas real-world data typically cannot.
@@ 4.1.1 Setting
We generate two classes of data by randomly and independently drawing the samples from two different random vectors $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{n}\}^{\top}$ and $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}^{\top}$, for which we set $X_{i}\sim N(\mu_{i},\sigma_{i}^{2})$ and $Y_{i}\sim N(-\mu_{i},\sigma_{i}^{2})$, with $\mu_{i}\in(-1,0)\cup(0,1)$ and $\sigma_{i}^{2}=1-\mu_{i}^{2}$, according to the data distributions specified in Property 1. Considering the fact that the features of real-world data usually exhibit sparse structures (Weiss and Freeman, 2007; Kotz et al., 2012), we further suppose that the means $\mu_{i}$ decay exponentially in magnitude, i.e. $|\mu_{i+1}|/|\mu_{i}|=e^{-\lambda}$, $\lambda\geq 0$, and set $\mu_{1}=0.8$ in the following simulation. It can be seen that with the increasing of $\lambda$, the mean’s magnitude $|\mu_{i}|$ (with $i>1$) will become smaller, indicating a smaller data element $X_{i}$ (in magnitude) and a sparser data structure. However, the data element $X_{i}$ with smaller $\mu_{i}$, is not favorable for quantization to improve feature discrimination, as indicated by previous numerical analyses. The impact of data sparsity on quantization can be investigated by increasing the value of the parameter $\lambda$.
With the data model described above, we randomly generate two classes of data, each class containing 1000 samples. The dataset is split into two parts for training and testing, in a ratio of 4:1. Then we evaluate the KNN and SVM classification on them. The classification accuracy is determined by averaging the accuracy results obtained from repeating the data generation and classification process 100 times. The major results for KNN with Euclidean distance are presented in the main text, as shown in Figures 2 and 3; and other results, such as those for KNN with cosine distance and SVM, are given in Appendix C, specifically in Figures 9 to 13. It can be seen that the three classifiers exhibit similar performance trends. For brevity, we will focus more on the results of KNN with Euclidean distance in the subsequent discussion.
@@ 4.1.2 Results
In Figure 2, we investigate the classification performance for the data generated with different parameters $\lambda\in\{0,0.01,0.1,1\}$, namely with different sparsity levels. Recall that the larger the $\lambda$, the smaller the $|\mu_{i}|$, or say the smaller the data element $X_{i}$ (in magnitude). By previous analyses, the data element $X_{i}$ with smaller $|\mu_{i}|$ is not conducive to enhancing feature discrimination through quantization. Nevertheless, empirically, the negative effect does not appear to be significant. From Figure 2, it can be seen that when increasing $\lambda$ from $0.1$ to $1$, there have been quantization thresholds $\tau$ that can yield better classification performance than original data. In addition, it noteworthy that as $\lambda$ increases, the overall classification accuracy of original data will decrease. This decreasing trend also impacts the absolute performance of the quantized data, even though it may outperform original data.
@@
The impact of data dimensions $n\in\{1,100,10000\}$ on classification is investigated in Figure 3, where the data are generated with the exponentially decaying parameter $\lambda=1$. It can be seen that with the increasing of data dimension, the range of the quantization thresholds $\tau$ that outperform original data tends to expand, but the performance advantage declines. As previously discussed, the decline should be attributed to the data element $X_{i}$ with small means $|\mu_{i}|$, whose quantity will rise with the data dimension $n$, particularly when the decay parameter $\lambda$ of $|\mu_{i}|$ is large. To alleviate this adverse effect, it is recommended to choose a relatively smaller $\lambda$ for high-dimensional data, indicating a structure that is not overly sparse. Conversely, when the high-dimensional data is highly sparse, we should reduce its dimension to improve the classification performance under quantization.
@@
Combining the results in Figures 2, 3, and 912, we can say that both KNN and SVM enable quantization to improve the classification accuracy of original data, within specific ranges of quantization thresholds $\tau$. If closely examining these ranges, it can be observed that KNN often performs better when using Euclidean distance than using cosine distance. This can be attributed to the advantage of Euclidean distance over cosine distance in measuring the distance between 00 and ±$\pm 1$. Also, KNN often outperforms SVM, such as the case of $\lambda=0.1$ shown in Figures 2 and 10. This is because the support vector of SVM relies on a few data points located on the boundary between two classes, which may deteriorate during quantization. In contrast, KNN depends on the high-quality data points within each class, making it resilient to quantization noise.
@@
In Figure 13, we present the KNN classification accuracy for the binary, ternary, and original synthetic data generated with the decay parameter $\lambda=1$ and data dimension $n=1$. For comparisons, we also provide the feature discrimination values and quantization errors across different thresholds $\tau$ for both binary and ternary data. It can be seen that higher classification accuracy tends to correlate with larger feature discrimination values, rather than smaller quantization errors. This suggests that the fluctuation in classification accuracy across different $\tau$ values are more appropriately reflected by feature discrimination, rather than quantization errors.

### v2

# Binary and Ternary Quantization Can Enhance Feature Discrimination

@@ Abstract
Quantization is widely applied in machine learning to reduce computational and storage costs for both data and models. Considering that classification tasks are fundamental to the field, it is crucial to investigate how quantization impacts classification performance. Traditional research has focused on quantization errors, assuming that larger errors generally lead to lower classification accuracy. However, this assumption lacks a solid theoretical foundation and often contradicts empirical observations. For example, despite introducing significant errors, $\{0,1\}$-binary and $\{0,\pm 1\}$-ternary quantized data have sometimes achieved classification accuracy comparable or even superior to full-precision data. To reasonably explain this phenomenon, a more accurate evaluation of classification performance is required. To achieve this, we propose a direct analysis of the feature discrimination of quantized data, instead of focusing on quantization errors. Our analysis reveals that both binary and ternary quantization can potentially enhance, rather than degrade, the feature discrimination of the original data. This finding is supported by classification experiments conducted on both synthetic and real data.
@@ 1 Introduction
For instance, it has been observed that some extremely low bit-width quantization methods, such as 1-bit $\{0,1\}$-binary and 2-bit $\{0,\pm 1\}$-ternary quantization, which have been successively applied in large-scale retrieval (Charikar, 2002) and deep network quantization (Qin et al., 2020; Gholami et al., 2022; Wang et al., 2023), can achieve comparable or even superior classification performance than their full-precision counterparts (Courbariaux et al., 2015; Lin et al., 2016b; Lu et al., 2023), despite suffering from high quantization errors. Apparently, the remarkable classification improvement resulting from quantization should not be attributed to the significant quantization errors. This reveals the inadequacy of quantization errors in evaluating the actual classification performance. Due to the absence of a theoretical explanation, the classification improvement induced by quantization has often been regarded as incidental and received little attention. Instead of quantization errors, in the paper we demonstrate that this intriguing phenomenon can be reasonably explained by feature discrimination. Following the Fisher’s linear discriminant analysis (Fisher, 1936), we here refer to feature discrimination as the ratio between inter-class and intra-class scatters, and evaluate the classification performance based on the rule that the higher the feature discrimination, the easier the classification. To the best of our knowledge, this is the first study that exploits feature discrimination to investigate the impact of quantization on classification, although it is more direct and reasonable than quantization errors in evaluating classification performance. The scarcity of relevant research may be attributed to the nonlinearity of the quantization operation, which substantially increases the analytical complexity of feature discrimination functions.
In the paper, it is demonstrated that the impact of the threshold-based binary and ternary quantization on feature discrimination can be analyzed, when the data are appropriately modeled using a Gaussian mixture model, with each Gaussian element representing one class of data. The Gaussian mixture model is chosen here based on two considerations. Firstly, the model has been well-established for approximating the distributions of real-world data (Torralba and Oliva, 2003; Weiss and Freeman, 2007) and their feature transformations (Wainwright and Simoncelli, 1999; Lam and Goodman, 2000). Secondly, the closure property of Gaussian distributions under linear operations can simplify the analysis of the feature discrimination function. By analyzing the discrimination across varying quantization thresholds, it is found that there exist certain quantization thresholds that can enhance the discrimination of original data, thereby yielding improved classification performance. This finding is validated through extensive classification experiments on both synthetic and real data, covering diverse modalities such as images, speech, and text.
@@ 2.1 Quantization functions
where the threshold parameter $\tau\in(-\infty,+\infty)$ for $f_{b}(x;\tau)$, and $\tau\in[0,+\infty)$ for $f_{t}(x;\tau)$. The two functions operate element-wise on a vector $\mathbf{x}=(x_{1},x_{2},\cdots,x_{n})^{\top}\in\mathbb{R}^{n}$, namely $f_{b}(\mathbf{x};\tau)=(f_{b}(x_{1};\tau),f_{b}(x_{2};\tau),\cdots,f_{b}(x_{n};\tau))^{\top}$ and the same applies to $f_{t}(\mathbf{x};\tau)$.
@@ 2.2 Data distributions
To facilitate data processing, we commonly standardize variables $X$ and $Y$ using Z-score standardization, which induces specific relationships between their distributions. More precisely, in a binary classification problem, the dataset we handle is a mixture, denoted as $Z$, comprising two classes of samples drawn respectively from $X$ and $Y$. Usually, the mixture $Z$ is assumed to possess a balanced class distribution, meaning that samples are drawn from $X$ and $Y$ with equal probabilities. Under this assumption, when we perform standardization by subtracting the mean and dividing by the standard deviation for each sample in $Z$, the distributions of $X$ and $Y$ (in $Z$) will become
@@
From Equations (3) and (4), it can be seen that after standardization, the two classes of variables $\tilde{X}$ and $\tilde{Y}$ still exhibit Gaussian distributions, but showcase two interesting properties: 1) their means are symmetric about zero; and 2) they have the sum of the square of the mean and the variance equal to one. In the paper, we will focus our study on such standardized data, which can be characterized as follows.
Property 1 (Distributions of two classes of standardized data)
Consider a dataset comprising two classes of data, each sampled equally from two Gaussian distributions with distinct means but identical variances. Upon standardization, the two classes of data will follow the distributions $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$, respectively, where the relationship $\mu^{2}+\sigma^{2}=1$ holds, with $\mu\in(0,1)$.
@@ 2.4 Goal
The major goal of the paper is to investigate whether there exist threshold values $\tau$ in the binary quantization $f_{b}(x;\tau)$ and the ternary quantization $f_{t}(x;\tau)$, such that the quantization can enhance the feature discrimination of original data, namely having $D_{b}>D$ and $D_{t}>D$.
@@ 3.1 Theoretical results
Remarks: Regarding the two theorems, there are three issues worth discussing. 1) The two theorems suggest that both binary and ternary quantization methods can indeed enhance the classification performance of original data, if there exist quantization thresholds $\tau$ that can satisfy the constraints shown in Equations (8) and (9). The following numerical analysis demonstrates that the desired threshold $\tau$ does exist, when the two classes of data $X\sim N(\mu,\sigma^{2})$ and $Y\sim N(-\mu,\sigma^{2})$ are assigned appropriate values for $\mu$ and $\sigma$. This threshold $\tau$ can be typically determined using gradient descent methods, as detailed in Appendix B. 2) Our theoretical analysis is based on the premise that the data vectors within the same class follow Gaussian distributions in each dimension of the vectors. This condition should hold true when two classes of data are readily separable, as in this case the data points within each class should cluster tightly, allowing for Gaussian approximation. This explains the recent research findings, that the binary or ternary quantization tends to achieve comparable or superior classification performance, when handling relatively simple datasets (Courbariaux et al., 2015; Lin et al., 2016b), or distinguishable features (Lu et al., 2023). 3) The conclusion we derive in Theorem 4 for $\{0,1\}$-binary quantization also applies to another popular $\{-1,1\}$-binary quantization (Qin et al., 2020), since the Euclidean distance of the former is equivalent to the cosine distance of the latter.
@@ 3.2 Numerical validation
In this part, we conduct numerical analyses for two primary objectives. Firstly, we aim to prove the existence of the desired quantization threshold $\tau$ that holds Equations (8) and (9), namely making the left sides of the two inequalities larger than their right sides (with values equal to zero). For this purpose, we compute the values of the left sides of Equations (8) and (9), through assigning specific values to $\tau$, as well as to the two variables $X$ and $Y$’s distribution parameters $\mu$ and $\sigma^{2}$. Note that we here set $\sigma^{2}=1-\mu^{2}$, $\mu\in(0,1)$, in accordance with Property 1. In Figure 1, we examine the case that fixes $\mu=0.8$ and $\sigma^{2}=0.36$, while varying the value of $\tau$ with a step width 0.01. The results for binary quantization and ternary quantization are provided in Figures 1 (a) and (c), respectively. It can be seen that for the two quantization methods, their conditions shown in Equations (8) and (9) will hold when respectively having $\tau\in[-0.2,0.2]$ and $\tau\in[0,0.5]$. This proves the existence of the desired quantization threshold $\tau$ that can enhance feature discrimination. For limited space, we here only discuss the case of $\mu=0.8$ (and $\sigma^{2}=1-\mu^{2}$) in Figure 1. By examining different $\mu\in(0,1)$ in the same way, we can find that the quantization threshold $\tau$ that holds Equations (8) and (9), is present when $\mu\in(0.76,1)$ and $\mu\in(0.66,1)$, respectively; see Figures 8 and 9 for more evidences. This result implies two consequences. On one hand, ternary quantization has more chances to enhance feature discrimination compared to binary quantization, as the former has a broader range of $\mu$. On the other hand, the enhanced discrimination tends to be achieved when $\mu$ is sufficiently large, coupled with a correspondingly small $\sigma$, or when the discrimination between two classes of data is sufficiently high.
@@
The second goal is to verify that the quantization thresholds $\tau$ we estimate with Equations (8) and (9) in Theorems 4 and 5, can indeed enhance feature discrimination. To this end, it needs to prove that the ranges of $\tau$ derived by Equations (8) and (9), such as the ones depicted in Figures 1 (a) and (c), are consistent with the ranges we can statistically estimate by the discrimination definitions $D$, $D_{b}$ and $D_{t}$, as specified in Definitions 2 and 3. To estimate the discrimination $D$, $D_{b}$ and $D_{t}$, we randomly generate 10,000 samples from $X\sim N(0.8,0.36)$ and $Y\sim N(-0.8,0.36)$, respectively, and then statistically estimate them with Equations (5), (6) and (7), across varying $\tau$ (with a step width 0.01). The results are provided in Figures 1 (b) and (d), respectively for binary and ternary quantization. It can be seen that we have $\tau\in[-0.2,0.2]$ for $D_{b}>D$, and have $\tau\in[0,0.5]$ for $D_{t}>D$. The results coincide with the theoretical results shown in Figures 1 (a) and (c), validating Theorems 4 and 5.
@@ 4 Experiments
Through previous theoretical and numerical analyses, we have demonstrated that binary and ternary quantization can enhance feature discrimination between two classes of data, when the data vectors within each class exhibit Gaussian distributions across each dimension of their feature vectors. As enhanced feature discrimination is expected to improve classification performance, this section aims to confirm this improvement by evaluating classification performance. To see the generalizability of our theoretical findings, we evaluate classification performance on both synthetic and real data. Synthetic data can accurately adhere to the distribution conditions specified in our theoretical analysis, whereas real-world data typically cannot.
@@ 4.1.1 Setting
We generate two classes of data by randomly and independently drawing the samples from two different random vectors $\mathbf{X}=\{X_{1},X_{2},\cdots,X_{n}\}^{\top}$ and $\mathbf{Y}=\{Y_{1},Y_{2},\cdots,Y_{n}\}^{\top}$, for which we set $X_{i}\sim N(\mu_{i},\sigma_{i}^{2})$ and $Y_{i}\sim N(-\mu_{i},\sigma_{i}^{2})$, with $\mu_{i}\in(-1,0)\cup(0,1)$ and $\sigma_{i}^{2}=1-\mu_{i}^{2}$, according to the data distributions specified in Property 1. Considering the fact that the features of real-world data usually exhibit sparse structures (Weiss and Freeman, 2007; Kotz et al., 2012), we further suppose that the means $\mu_{i}$ decay exponentially in magnitude, i.e. $|\mu_{i+1}|/|\mu_{i}|=e^{-\lambda}$, $\lambda\geq 0$, and set $\mu_{1}=0.8$ in the following simulation. It can be seen that with the increasing of $\lambda$, the mean’s magnitude $|\mu_{i}|$ (with $i>1$) will become smaller, indicating a smaller data element $X_{i}$ (in magnitude) and a sparser data structure. However, the data element $X_{i}$ with smaller $\mu_{i}$, is not favorable for quantization to enhance feature discrimination, as indicated by previous numerical analyses. The impact of data sparsity on quantization can be investigated by increasing the value of the parameter $\lambda$.
With the data model described above, we randomly generate two classes of data, each class containing 1000 samples. The dataset is split into two parts for training and testing, in a ratio of 4:1. Then we evaluate the KNN and SVM classification on them. The classification accuracy is determined by averaging the accuracy results obtained from repeating the data generation and classification process 100 times. The major results for KNN with Euclidean distance are presented in the main text, as shown in Figures 2 to 4; and other results, such as those for KNN with cosine distance and SVM, are given in Appendix C, specifically in Figures 10 to 13. It can be seen that the three classifiers exhibit similar performance trends. For brevity, we will focus more on the results of KNN with Euclidean distance in the subsequent discussion.
@@ 4.1.2 Results
In Figure 2, we investigate the classification performance for the data generated with different parameters $\lambda\in\{0,0.01,0.1,1\}$, namely with different sparsity levels. Recall that the larger the $\lambda$, the smaller the $|\mu_{i}|$, or say the smaller the data element $X_{i}$ (in magnitude). By previous analyses, the data element $X_{i}$ with smaller $|\mu_{i}|$ is not conducive to enhancing feature discrimination through quantization. Nevertheless, empirically, the negative effect does not appear to be significant. From Figure 2, it can be seen that when increasing $\lambda$ from $0.1$ to $1$, there have been quantization thresholds $\tau$ that can yield better classification performance than the original, full-precision data. In addition, it noteworthy that as $\lambda$ increases, the overall classification accuracy of original data will decrease. This decreasing trend also impacts the absolute performance of the quantized data, even though it may outperform original data.
@@
The impact of data dimensions $n\in\{1,100,10000\}$ on classification is investigated in Figure 3, where the data are generated with the exponentially decaying parameter $\lambda=1$. It can be seen that with the increasing of data dimension, the range of the quantization thresholds $\tau$ that outperform original data tends to expand, but the performance advantage declines. As previously discussed, the decline should be attributed to the data element $X_{i}$ with small means $|\mu_{i}|$, whose quantity will rise with the data dimension $n$, particularly when the decay parameter $\lambda$ of $|\mu_{i}|$ is large. To alleviate this adverse effect, it is recommended to choose a relatively smaller $\lambda$ for high-dimensional data, indicating a structure that is not overly sparse. Conversely, when the high-dimensional data is highly sparse, we should reduce its dimension to enhance the classification performance under quantization.
@@
Combining the results in Figures 2, 3, and 1013, we can say that both KNN and SVM enable quantization to improve the classification accuracy of original data, within specific ranges of quantization thresholds $\tau$. If closely examining these ranges, it can be observed that KNN often performs better when using Euclidean distance than using cosine distance. This can be attributed to the advantage of Euclidean distance over cosine distance in measuring the distance between 00 and ±$\pm 1$. Also, KNN often outperforms SVM, such as the case of $\lambda=0.1$ shown in Figures 2 and 11. This is because the support vector of SVM relies on a few data points located on the boundary between two classes, which may deteriorate during quantization. In contrast, KNN depends on the high-quality data points within each class, making it resilient to quantization noise.
(a) Classification accuracy (B)
(b) Discrimination value (B)
(c) Quantization error (B)
(d) Classification accuracy (T)
(e) Discrimination value (T)
(f) Quantization error (T)
Figure 4: KNN (Euclidean distance) classification accuracy, feature discrimination values and quantization errors are provided for the synthetic data generated with the decay parameter $\lambda=1$ and data dimension $n=1$, as well as for their binary and ternary quantization derived across different thresholds $\tau$.
@@
As theoretically predicted, Figure 4 illustrates that higher classification accuracy tends to be achieved at quantization thresholds $\tau$ that result in greater feature discrimination values, rather than smaller quantization errors. In particular, the variation in classification accuracy with respect to $\tau$ is more closely associated with feature discrimination values than with quantization errors. This confirms the superiority of our feature discrimination analysis in assessing the classification performance of quantized data, compared to the traditional approach of using quantization errors.